OpenStudy (anonymous):

Solve the following system of equations. 2x – y + z = –3 2x + 2y + 3z = 2 3x – 3y – z = –4 (1, 3, 2) (–1, 3, 2) (1, –3, 2) (1, 3, –2)

OpenStudy (lgbasallote):

if i were you i'll add the first equation and third equation first...

OpenStudy (anonymous):

if you've got the choices,check which triplet satisfies the 3 equations

OpenStudy (lgbasallote):

that's too long of a process :C and no fun too >:))

OpenStudy (anonymous):

So you have a couple choices. You can check the possible answers. If you do that, i'd recommend using the second equation first.. You can do elimination. You can use Cramer's rule. 2 -1 1 | -3 2 2 3 | 2 3 -3 -1 | 4 Let's do elimination.

OpenStudy (anonymous):

taking too long, try eliminating z first I add eqn 1 to eqn 3, also -3 times eqn 1 plus eqn 2, I get 5x - 4y = -7 -4x + 5y = 11 Peeking at the answers, x = 1, y = 3 works. Try in eqn 1 I get z = -2 Check my work, I could have made a mistake!